OpenStudy (anonymous):

(Theory number) Find all prime numbers p, such that 7p = 8x^2 - 1 P^2 = 2y^2 - 1 with x, y are integers.

OpenStudy (anonymous):

By using trials i got p = 41. But im not sure if many values other of p.

ganeshie8 (ganeshie8):

7p = 8x^2 - 1 p^2 = 2y^2 - 1 subtract and get p(p-7) = 2(y-2x)(y+2x)

OpenStudy (anonymous):

Ok. So, what next ?

ganeshie8 (ganeshie8):

not sure, still working on it..

OpenStudy (anonymous):

7p = 8x^2 - 1 p^2 = 2y^2 - 1 according the first eq, we know that p is odd prime. Subtract, get p(p-7) = 2(y-2x)(y+2x) p-7 = even = 2(y - 2x) = 2y - 4x and p = y + 2x subtract again, get -7 = y - 6x or y = 6x - 7 here we can see the relation betwen x and y :)

OpenStudy (anonymous):

Ah, yes. Just subtitute y = 6x - 7 into p^2 = 2y^2 - 1, it become P^2 = 2(6x - 7)^2 - 1 p^2 = 72x^2 - 84x + 97 From the equation., 7p = 8x^2 - 1 Square both side, get 49p^2 = 64x^4 - 16x^2 + 1 49(72x^2 - 84x + 97) = 64x^4 - 16x^2 + 1 then solve for x ...

OpenStudy (anonymous):

Ask wolfram only x = 1 and x = 6, as integers. Check one by one only x = 6 causes p be a prime. http://m.wolframalpha.com/input/?i=64x%5E4+-+3544x%5E2+%2B+8232x+-+4752+%3D+0&x=0&y=0